OpenStudy (anonymous):

Vickie pays $90.00 for a new dress at a President's Day sale. If the original price for the dress was $225.00, what is the percent discount?

OpenStudy (anonymous):

let x be the % discount... implying x% of 225 = 90... which implies (x/100)*225=90... implying 2.25*x=90... implying x= 90/2.25 implying.. x=40.. so its a 40% dicount.

ganeshie8 (ganeshie8):

\(\huge \frac{225-90}{225}*100 \\ \huge = \frac{35}{225}*100 \\ \huge = ?\)

OpenStudy (anonymous):

It's simply the marked price- selling price/ marked price x 100 225-90/225 x 100 135/225 x100= 60%

OpenStudy (anonymous):

Is 60% answer